OpenStudy (anonymous):

solve for x: log_3(2x-1)=2

OpenStudy (anonymous):

this is equivalent to 3^2=2x-1 2x=10 x=5 I assumed you meant log base 3 of 2x-1=2

OpenStudy (anonymous):

5

OpenStudy (anonymous):

yes it did thats how we write it on out internet homework

OpenStudy (anonymous):

ok then 5 is correct

OpenStudy (anonymous):

2x-1=9 therefore x=5